Solve : x(y^2-z^2)p+y(z^2-x^2)q=z(x^2-y^2)
1
Expert's answer
2020-05-04T18:32:50-0400

The given equation is a Lagrange's linear equation. The auxiliary equations are


"\dfrac{dx}{x(y^{2}-z^{2})}=\dfrac{dy}{y(z^{2}-x^{2})}=\dfrac{dz}{z(x^{2}-y^{2})}" .


Using the multipliers, "x,y,z" each of the equation is equal to

"\dfrac{xdx+ydy+zdz}{x^{2}(y^{2}-z^{2})+y^{2}(z^{2}-x^{2})+z^{2}(x^{2}-y^{2})}"


Therefore,

"xdx+ydy+zdz=0\\

\text{Integrating, we get}\\

\dfrac{x^{2}}{2}+\dfrac{y^{2}}{2}+\dfrac{z^{2}}{2}=c\\
x^{2}+y^{2}+z^{2}=c_{1}~~~~~~~~~~~~(1)"

Now, using the multipliers, "\dfrac{1}{x}, \dfrac{1}{y}, \dfrac{1}{z}" each of the equation is equal to

"\dfrac{\frac{1}{x}dx+\frac{1}{y}dy+\frac{1}{z}dz}{y^{2}-z^{2}+z^{2}-x^{2}+x^{2}-y^{2}}"


Therefore,

"\frac{1}{x}dx+\frac{1}{y}dy+\frac{1}{z}dz=0\\
\text{Integrating, we get}\\
\log x+\log y+\log z = \log c_{2}\\
\log (xyz)=\log c_{2}\\
xyz = c_{2}~~~~~~~~~~~~~~~~~~~~~~~~(2)"


The general solution is given by,

"\phi(c_{1},c_{2})=0\\
\phi(x^{2}+y^{2}+z^{2},xyz)=0\\"